OpenStudy (anonymous):

the sum of two angles is 180. one angle measures 3a and the other angle measures 2a-25. find a. then find the measure of each angle.

OpenStudy (anonymous):

3a+2a-25=180. 5a-25=180 5a=205 a=41

OpenStudy (anonymous):

what about the measure of each angle?

OpenStudy (anonymous):

One angle is 3a=123 and the other is 2a-25=57

OpenStudy (anonymous):

thanks